A sample of gas contains 0.2000 mol of CH(g).mil 0
Answer #1

CH4 + H2O à 3 H2 + CO

0.2 mol of methane and 0.2 mol of H2O gives 3* 0.2 mol of H2 and 1* 0.2mol of CO

Total mole of reactants n1 = 0.4 mol

Total mol of products n2 = 0.8 mol

Avagadro’s law:

V1/n1 = V2/n2 (at constant pressure and temperature)

15.9 L/0.4 m = V2/0.8 m

V2 = 31.8 L